OpenStudy (anonymous):

r squared+ 9 squared=(6+r)squared

OpenStudy (dtan5457):

\[r^2+9^2=(6+r)^2\]

OpenStudy (dtan5457):

Correct?

OpenStudy (dtan5457):

Did I type the question correctly?

OpenStudy (dtan5457):

Well, your not responding. So, I'm gonna continue assuming that is the question.

OpenStudy (dtan5457):

Further simplify the expression to \[r^2+81=36+12r+r^2\]

OpenStudy (dtan5457):

Cross out the r^2 as it is in both sides. You are left with 81=36+12r Subtract 36. 45=12r Solve for r now.

OpenStudy (anonymous):

thats it i lost connection how did you get 12r

OpenStudy (anonymous):

r would equal 3.75

OpenStudy (dtan5457):

3.75 is correct.

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

Hm?

OpenStudy (anonymous):

how did you get 12r

OpenStudy (dtan5457):

(6+r)(6+r) You foil that.

OpenStudy (dtan5457):

You end up with 36+12r+r^2

OpenStudy (anonymous):

oh! i remember now thanks you i am in geometry and totally forgot how to do that @dtan5457 thank you

OpenStudy (dtan5457):

No problem. :)